Consider the numbers 1,2,3 and 5. If the same number is added to each of the numbers, it is found that the ratio of the first new number to the second is the same as the ratio of the third new number to the fourth. Find the number.
Answer by edjones(8007)

Let x be the number that is added.
(x+1)/(x+2)=(x+3)/(x+5)
(x+1)(x+5)=(x+3)(x+2)
x^2+6x+5=x^2+5x+6
x=1
.
Ed